# A Survey on Adversarial Contention Resolution

### v3

@@ 2.2. Deterministic and Hybrid Approaches
In contrast to randomized backoff, there are a handful of results that propose deterministic approaches for what is, essentially, the case of static-arrivals. Specifically, packets may arrive over time. However, once a collision occurs, those participants whose packets collided must execute a “conflict resolution” algorithm, while all other participants learn about the collision via channel feedback and refrain from transmitting; that is, no other packets are sent until this conflict is resolved.
For these results, there are $k\geq 2$ participants, each with a unique ID in the range $\{1,...,n\}$. Komlos and Greenberg (komlos:asymptotically, ) show the existence of a non-adaptive, deterministic algorithm that, with knowledge of $k$ and $n$, has a makespan of $O(k+k\log(n/k))$; their approach requires that each participant $u$ only adapts to channel feedback during slots in which $u$ transmits. Related to this, in the static case, Kowalski (DBLP:conf/podc/Kowalski05, ) proves the existence of a non-adaptive deterministic algorithm with $O(k\log(n/k))$ makespan, and also gives an explicit construction for a non-adaptive algorithm with $O(k\,\texttt{polylog}(n))$ makespan. These results stand in contrast to prior work by Capetanakis (capetanakis:generalized, ) and Tsybakov and Mikhailov (tsybakov1978free-russian, ; tsybakov1978free, ) who give deterministic algorithms with the same asymptotic makespan without needing to know $k$; however, their algorithms require that stations receive feedback even in slots where they themselves are not transmitting. Related to these results, Greenberg, Flajolet, and Ladner (DBLP:journals/jacm/GreenbergFL87, ; DBLP:conf/focs/GreenbergL83, ) give a randomized algorithm to estimate the number of stations that need to send a packet; this relies on collision detection. This is combined with prior deterministic approaches (capetanakis:generalized, ; hayes:adaptive, ; tsybakov1978free-russian, ; tsybakov1978free, ; fayolle1982capacity, ; hofri1984stack, ), yielding several hybrid algorithms, and the best expected makespan is at most $2.1338n+o(n)$.
@@
We introduce the wakeup problem, which is described as follows. There are $n$ participants, of which some are active (or “awake”), and the remainder are inactive (or“sleeping”) with only the ability to listen. Each awake participant has a packet to send, and upon the first success, all sleeping participants “wake up” and the problem is solved. Typically, the metric of interest is the number of rounds until the first success. Thus, contention resolution differs from wakeup: it requires all participants to succeed, while wakeup hinges on just the first success.
@@
Despite this difference, historically the wakeup problem has considerable overlap in the literature with contention resolution, and is sometimes referred to by the same name (e.g., (fineman:contention, ; fineman:contention2, ; gilbert2021contention, ; DEMARCO20171, ; de2013contention, )). To continue the naming complications, in the static case where the set of active participants are fixed from the start, the wakeup problem is sometimes referred to as leader election (LE).
@@
Static wakeup and LE may be viewed as a natural starting point for solving contention resolution, as repeatedly applying them over all $n$ packets offers a solution, albeit an inefficient one. A more meaningful connection can be drawn between static wakeup and contention resolution: wake-up protocols often produce an estimate for the $n$, the size of the system, which may be used to bootstrap a contention resolution protocol. More specifically, many algorithms are “uniform” or ‘fair”: in any fixed round, each participant sends packets with the same probability (independently of the prior communication history). For such algorithms, the reciprocal of the sending probability when the first success occurs may yield an accurate estimate of $n$; obtaining system-size knowledge is a well-known component of contention resolution (e.g., (jurdzinski:energy, ; chang:exponential-jacm, ; gilbert2021contention, )). For instance, such an estimate can be used to set the first window size in algorithms such as FB and STB/TSTB (recall Section 2.1), bypassing much of the backing off that would otherwise be required, and speeding up the runtime by a constant factor.
@@
For static wakeup, makespan refers to number of slots until the first success (recall Section 1.1.3). An early result on the wakeup problem is by Willard (willard:loglog, ), who considers the case with CD. Here, Willard establishes a lower bound on uniform algorithms—for any fixed round, each participant sends packets with the same probability (independently of the prior communication history)—showing that the expected makespan is at least $\lfloor\lg\lg n\rfloor-O(1)$, and an upper bound of $\lg\lg n+O(1)$ for expected makespan is also given.
@@
In the case of no-CD and where $n$ is known, Kushilevitz and Mansour (kushilevitz1998omega, ; kushilevitz1993omega, ) prove a lower bound of $\Omega(\log n)$ expected makespan; this applies even to unfair algorithms. An asymptotically-matching upper bound on expected makespan, and a bound of $O(\log^{2}n)$ w.h.p. is achieved by Bar-Yehuda et al. (bar1992time, ). In subsequent work that unifies many prior lower-bound arguments, Newport (newport2014radio, ; newport:radio-journal, ) also shows $\Omega(\log n)$ expected makespan and $\Omega(\log^{2}n)$ makespan w.h.p. for any randomized algorithms; these bounds hold even with multiple channels and/or CD. For algorithms that succeed with probability at least $1-\epsilon$, this lower bound was strengthened to $\Omega(\log n\,\log(1/\epsilon))$ by Farach-Colton et al. (farach-colton:lower-bounds, ). We note a couple things. First, these lower bounds are dramatically different from the lower bound by Willard (willard:loglog, ), where CD is available. Second, Jurdziński and Stachowiak (jurdzinski2002probabilistic, ) provide an algorithm with $O(\log n\,\log(1/\epsilon))$ makespan so long as at least one of the following is true: devices have unique labels, or $n$ is known.
@@
The signal-to-interference-plus-noise ratio (SINR) model has been considered in the context of several fundamental wireless-communication problems (for details, see (moscibroda2006complexity, ; jurdzinski2012distributed, ; jurdzinski2013distributed, ; moses2020deterministic, ; jurdzinski2015cost, ; JurdzinskiR17, )). Informally, under the SINR model, the ratio of (i) the power of an incoming transmission to (ii) the sum interference from other devices plus the floor noise, must exceed a threshold value in order for the transmission to succeed. In this setting, Fineman et al. (fineman:contention, ) shows how to solve wakeup with $O(\log n+\log R)$ makespan w.h.p., where $R$ is ratio between the longest and shortest link. The authors point out that, in many practical settings, $R$ is at most a polynomial in $n$, which yields an $O(\log n)$ result. [^5] This ratio appears in other related results under the SINR model, such as interference scheduling or link scheduling (e.g., (fanghanel2011improved, )). The authors also show a general lower bound (e.g., it is not restricted to the class of fair algorithms) of $\Omega(\log n)$.
Finally, we note a recent result by Gilbert et al. (gilbert2021contention, ), that examines the implications of system-size “predictions”; that is, the algorithm is given access to the probability distribution defined over the possible values of $n$. The authors give lower and upper bounds for both deterministic and randomized solutions, for the CD and no-CD cases. Interestingly, in addition to new bounds, this work recovers some lower bounds, including the $\Omega(\log\log n)$ lower bound by Willard (willard:loglog, ) mentioned above.
Summary. For static contention resolution, many randomized algorithms have been proposed, with STB providing one of the strongest with-high-probability results: asymptotically-optimal makespan with no-CD, no knowledge of $n$, and only $O(\log n)$ channel-access attempts.
In contrast, purely deterministic approaches cannot achieve the asymptotically-optimal makespan without some additional assumptions regarding CD, knowledge of $n$, or a unique labeling of participants. We see a similar discrepancy when we consider adversarial dynamic arrivals, which we explore next.
@@
For static wakeup, under each of CD or no-CD, asymptotically matching upper and lower bounds are known. It is interesting to compare these results to those of static contention resolution. In particular, while the makespan for contention resolution is higher, the optimal average per-packet time to success is $\Theta(1)$, which is lower than even the CD case for wakeup. As we will see, this is also a common theme in the dynamic case with randomized approaches.
@@ 3. Dynamic Arrivals
To this end, adversarial queuing theory (AQT) (borodin2001adversarial, ; cruz-one, ; cruz-two, ; andrews2001universal, ) has been used to model packet arrivals times, and many contention-resolution results fit within the AQT framework; that is, they address adversarial models that are parameterized by rates of packet injection and burstiness. Here, a key metric is stability, which we discuss in Section 3.1.
Another large body of the contention-resolution literature eschews the parameterization of AQT and instead considers a completely arbitrary schedule for when packets arrive. Here, we treat separately the literature on deterministic and randomized approaches, in Sections 3.2 and 3.3, respectively.
@@ 3.1. AQT and Contention Resolution
Several adversarial models appear in the AQT literature; see Borodin et al (borodin2001adversarial, ) for an in-depth discussion. In the context of contention resolution, there are two prominent models. The leaky-bucket adversary is defined such that, for any $t$ consecutive slots, at most $\rho t+b$ packets may be placed in the queue of any participant (cruz-one, ). For a contiguous segment of rounds, and any positive integer $w$, a window adversary of type $(\rho,w)$ can inject $\rho w$ packets in each contiguous segment of $w$ rounds into any set of stations; $w$ is referred to as a window.
@@
We can now describe why BEB’ is unstable for an arrival rate of $\Omega(\ln\ln n/\ln n)$; Figure 4 complements our discussion. Consider a window $W$ of $T=\Theta(n\ln n/\ln\ln n)$ slots. In the first slot, a single burst of $n/2$ packets arrives. This drives up the contention to $\Omega(n)$ in this slot. None of these packets is likely to succeed over the next $n/\ln n$ slots; to see this, note that in the $i$-th slot measured from this first slot is sending with probability $1/i$, yielding a contention of $\Omega(\ln n)$ in each of these $n/\ln n$ slots, and so the probability of success is $O(1/{\texttt{poly}({\ln n})})$ by Equation 1.
@@
However, given the burst of packets, the packets in the stream face $n/\ln n$ slots where contention is $\Omega(\ln n)$. Consequently, rather than succeeding, these packets accumulate in the system and, over these $n/\ln n$ slots, their contribution to contention is $\Omega(\ln\ln n/\ln n)\sum_{i=1}^{n/\ln n}1/i=\Omega(\ln\ln n)$ by the sum of a harmonic series. By Equation 1, the probability of success is $O(1/{\texttt{poly}({\ln n})})$, and thus the expected number of slots for a packet to succeed is $\Omega({\texttt{poly}({\ln n})})$.
@@
The remaining approaches in the literature are deterministic. Chlebus et al. (chlebus2009maximum, ; chlebus2007stability, ) consider adversaries where $\rho=1$; that is, one packet is injected per slot. The main positive result is a deterministic, adaptive protocol that guarantees against leaky-bucket adversaries that $O(n^{2}+b)$ packets are queued in the system at any given time, where $n$ is the number of participants. Here, $n$ is known to all participants and each ID is unique in the range $[1,...,n]$; this allows participants to schedule via passing a token, such that no collisions occur (notably, the positive results in this paper do not require collision detection). This result is shown to be asymptotically optimal in the sense that a leaky-bucket adversary with burstiness $2$ can force $\Omega(n^{2})$ total packets to be queued.
@@
Several impossibility results are also given—holding regardless of whether or not collision detection is available—that illustrate how even minor burstiness can prevent stability. Specifically, no full-sensing protocol can be stable (i) for at least two participants against a leaky-bucket adversary with burstiness $2$, (ii) or for at least three participants against a window adversary of burstiness $2$; and (iii) no acknowledgment-based algorithm is stable in a system of at least two participants against the window adversary of burstiness $2$.
@@
Additional impossibility results are made in the context of three natural properties. First, an algorithm is fair if each packet is transmitted successfully within $O(b/\rho)$ slots from the time it is injected. [^7] Note that this is a different notion of fairness than described in Section 2.3. The authors show that no algorithm can be fair and stable for any number of at least two participants against a leaky-bucket adversary of burstiness $2$. Second, an algorithm is withholding the channel if, upon successfully transmitting a packet, the station keeps sending until all of its pending packets are exhausted. No algorithm that withholds the channel can be stable for three stations against a leaky-bucket adversary of burstiness $2$. Third, an algorithm is said to be retaining if after a participant successfully transmits a packet, and if any packets arriving subsequent to this event are injected only at other stations, the participant eventually empties its queue. No retaining protocol is stable in a system of at least four stations against a window adversary of burstiness $2$. Chlebus et al. show that this implies no protocol that is fair can also be stable in a system of at least four stations against the window adversary of burstiness $2$.
@@
In contrast to the above, Chlebus et al. (chlebus2006adversarial, ; chlebus:adversarial, ) study deterministic algorithms under the same model as Chlebus et al. (chlebus2009maximum, ; chlebus2007stability, ), but where $\rho<1$. They show that strong stability is impossible for a rate $\omega(1/\log n)$, with or without collision detection. When collision detection is available, the authors give a full-sensing algorithm that is fair and stable for $\rho=O(1/\log n)$, and show that fairness is not possible for rates that are $\omega(1/\log n)$. When collision detection is not available, the authors give an existence proof for a stable and fair algorithm under a much smaller rate of $O(1/n\log^{2}n)$, and they give an explicit algorithm that achieves these properties for a rate of $O(1/n^{2}\log n)$. Finally, it is shown that no acknowledgement-based algorithm can be stable for a rate exceeding $3/(1+\lg n)$.
@@
Anantharamu et al. (anantharamu2010deterministic, ) and Aldawsari et al. (aldawsari:adversarial, ) design and analyze several deterministic algorithms that employ either a round-robin or token-passing approach for obtaining access to the channel. The authors derive bounds on latency under the leaky-bucket adversary model as a function of $\rho$, for $\rho<1/(2\log n)$ and general $b$. Note that this restriction on $\rho$ should not surprise us, given the negative results on strong stability by Chlebus et al. (chlebus2006adversarial, ; chlebus:adversarial, ) and prior results on the challenges posed by burstiness (chlebus2009maximum, ; chlebus2007stability, ). These algorithms are compared, both analytically and via simulations, against the randomized algorithms, BEB and Quadratic Backoff (QB), where the latter has each successive window size increase as the square of the previous window. The simulation results illustrate the impact of different injection rates, burstiness values, and the number of participants. These findings generally align with the theoretical predictions and, interestingly, the deterministic algorithms compare favorably to BEB and QB, suggesting that randomized approaches may not necessarily be superior. More-recent work by Hradovich et al. (HradovichKK21, ) provides a new average-case analysis of the maximum queue size and latency for several of the same algorithms.
@@
While many deterministic approaches require nodes to have unique IDs, Anantharamu and Chlebus (anantharamu2015broadcasting, ) consider deterministic algorithms where this is not true; furthermore, $n$ is unknown to the algorithm and may not even be finite. In trade, they focus on positive results against leaky bucket adversaries that have the additional constraint that they inject at most one packet per round. An algorithm is activation-based if stations that do not have any packets to send ignore any channel feedback. In terms of positive results, the authors provide a non-adaptive, activation-based algorithm that has bounded latency and is stable for an injection rate $\rho<1/3$ with CD, a non-adaptive full-sensing algorithm that has bounded latency and is stable for $\rho\leq 3/8$ with CD, and an adaptive activation-based algorithm that has bounded latency and is stable for injection rates less than $\rho\leq 1/2$ with no-CD.
@@
In terms of negative results, Anantharamu and Chlebus show that no deterministic algorithm can be fair against an adversary that is able to activate at most $2$ stations in any round—called a $2$-activating adversary—with $\lfloor\rho+b\rfloor=1+b\geq 2$. Additionally, the authors show that no deterministic acknowledgment-based algorithm is fair against a $1$-activating adversary for $\rho<1$ and $b\geq 3-2\rho$. This stands in contrast to the positive results in Chlebus et al. (chlebus2006adversarial, ; chlebus:adversarial, ) for $\rho<1$. Finally, Anantharamu and Chlebus prove that no deterministic algorithm with CD can provide bounded packet latency against a $1$-activating adversary with injection rate at least $3/4$ and $b\geq 1$. It is interesting to compare this result against the algorithm by Chlebus et al. (chlebus2009maximum, ; chlebus2007stability, ) discussed above, which achieves stability, having $O(n^{2}+b)$ total packets queued in the system at any given time (but where there are unique IDs and the number of stations is known).
@@
Anantharamu et al. (anantharamu2009adversarial, ; anantharamu2017adversarial, ) consider deterministic full-sensing and acknowledgment-based algorithms for the case where each participant is subject to an individual injection rate under a window adversary, where the window $w$ is unknown. There are $n$ participants, each with a unique ID in ${1,...,n}$, and $n$ is known. The authors prove a lower bound on latency of $\Omega(w\max\{1,\log_{w}n\})$ for $w\leq n$, and $\Omega(w+n)$ for $w>n$. Upper bounds are established by a full-sensing algorithm with CD, and by an adaptive algorithm without CD, each of which achieve a packet latency of $O(\min(n+w,w\log n))$; Anantharamu et al. also give a full-sensing algorithm without CD that has $O(n+w)$ queue size and $O(nw)$ latency.
In contrast to the above, the remaining results address a model where participants have unique IDs in $\{1,...,n\}$, where $n$ is known. Additionally, participants have access to a global clock. Hradovich et al. (hradovich2020contention, ) consider a problem variant where, in any fixed round, at most $k$ stations may access the multiple access channel; this is called a $k$-restrained channel. The authors design and analyze deterministic algorithms for leaky bucket adversaries with $\rho\leq 1$. The first is a stable adaptive algorithm whose maximum number of queued packets is $O(n^{2}+b)$, which is reminiscent of the bound in (chlebus2009maximum, ). The second is a full-sensing algorithm that is stable for $\rho\leq 1-1/n$. Finally, the authors give an acknowledgment-based algorithm with throughput $\Theta(k/n(\log^{2}n))$, which is complemented by a lower bound that no acknowledgement-based algorithm can achieve a throughput better than $\min\{k/n,1/(3\log n)\}$. Simulations indicate that the proposed algorithms perform favorably, with respect to queue size and number of channel accesses, relative to well-known backoff algorithms.

### v4

@@ 2.2. Deterministic and Hybrid Approaches
In contrast to randomized backoff, there are a handful of results that propose deterministic approaches for what is, essentially, the case of static-arrivals. Specifically, packets may arrive over time. However, once a collision occurs, those participants whose packets collided must execute a “conflict resolution” algorithm, while all other participants learn about the collision via channel feedback and refrain from transmitting until the conflict is resolved.
For these results, there are $k\geq 2$ participants, each with a unique ID in the range $\{1,...,n\}$. Komlós and Greenberg (komlos:asymptotically, ) show the existence of a non-adaptive, deterministic algorithm that, with knowledge of $k$ and $n$, has a makespan of $O(k+k\log(n/k))$; their approach requires that each participant $u$ only adapts to channel feedback during slots in which $u$ transmits. Related to this, Kowalski (DBLP:conf/podc/Kowalski05, ) proves the existence of a non-adaptive deterministic algorithm with $O(k\log(n/k))$ makespan; this is comparable to the lower bound by Greenberg and Winograd (greenberg1985lower, ) of $\Omega(k\log(n)/\log k)$, which applies to adaptive algorithms with collision detection.
Kowalski (DBLP:conf/podc/Kowalski05, ) also gives an explicit construction for a non-adaptive algorithm with $O(k\,\texttt{polylog}(n))$ makespan. These upper-bound results stand in contrast to prior work by Capetanakis (capetanakis:generalized, ) and Tsybakov and Mikhailov (tsybakov1978free-russian, ; tsybakov1978free, ) who give deterministic algorithms with the same asymptotic makespan without needing to know $k$; however, their algorithms require that stations receive feedback even in slots where they themselves are not transmitting. Related to these results, Greenberg, Flajolet, and Ladner (DBLP:journals/jacm/GreenbergFL87, ; DBLP:conf/focs/GreenbergL83, ) give a randomized algorithm to estimate the number of stations, which relies on collision detection; this is combined with prior deterministic approaches (capetanakis:generalized, ; hayes:adaptive, ; tsybakov1978free-russian, ; tsybakov1978free, ; fayolle1982capacity, ; hofri1984stack, ), yielding several hybrid algorithms, and the best expected makespan is at most $2.1338n+o(n)$.
@@
We introduce the wakeup problem, which is described as follows. There are $n$ participants, of which some are active (or “awake”), and the remainder are inactive (or“sleeping”) with only the ability to listen. Each awake participant has a packet to send, and upon the first success, all sleeping participants “wake up” and the problem is solved. Typically, the metric of interest is the number of rounds until the first success. Thus, contention resolution differs from wakeup: the former requires all participants to succeed, while the latter hinges on just the first success.
@@
Despite this difference, historically the wakeup problem has considerable overlap in the literature with contention resolution, and it is sometimes referred to by the same name (e.g., (fineman:contention, ; fineman:contention2, ; gilbert2021contention, ; DEMARCO20171, ; de2013contention, )). To continue the naming complications, in the static case where the set of active participants are fixed from the start, the wakeup problem is sometimes referred to as leader election (LE).
@@
Static wakeup and LE may be viewed as a natural starting point for solving contention resolution, as repeatedly applying them over all $n$ packets offers a solution, albeit an inefficient one. A more meaningful connection can be drawn between static wakeup and contention resolution: wake-up protocols often produce an estimate of $n$, which may be used to bootstrap a contention-resolution protocol. More specifically, many algorithms are “uniform” or ‘fair”: in any fixed round, each participant sends packets with the same probability (independently of the prior communication history). For such algorithms, the reciprocal of the sending probability when the first success occurs may yield an accurate estimate of $n$; obtaining system-size knowledge is a well-known component of solutions to contention resolution (e.g., (jurdzinski:energy, ; chang:exponential-jacm, ; gilbert2021contention, )). For instance, such an estimate can be used to set the first window size in algorithms such as FB and STB/TSTB (recall Section 2.1), bypassing much of the backing off that would otherwise be required, and speeding up the runtime by a constant factor.
@@
For static wakeup, the makespan refers to the number of slots until the first success (recall Section 1.1.3). An early result on the wakeup problem is by Willard (willard:loglog, ), who considers the case with CD. Here, Willard establishes a lower bound on uniform algorithms—for any fixed round, each participant sends packets with the same probability (independently of the prior communication history)—showing that the expected makespan is at least $\lfloor\lg\lg n\rfloor-O(1)$, and an upper bound of $\lg\lg n+O(1)$ for expected makespan is also given.
@@
In the case of no-CD and where $n$ is known, Kushilevitz and Mansour (kushilevitz1998omega, ; kushilevitz1993omega, ) prove a lower bound of $\Omega(\log n)$ expected makespan; this applies even to unfair algorithms. An asymptotically-matching upper bound on expected makespan, and a bound of $O(\log^{2}n)$ w.h.p. is achieved by Bar-Yehuda et al. (bar1992time, ). In subsequent work that unifies many prior lower-bound arguments, Newport (newport2014radio, ; newport:radio-journal, ) also shows $\Omega(\log n)$ expected makespan, and $\Omega(\log^{2}n)$ makespan w.h.p. for any randomized algorithms; these bounds hold even with multiple channels and/or CD. For algorithms that succeed with probability at least $1-\epsilon$, this lower bound was strengthened to $\Omega(\log n\,\log(1/\epsilon))$ by Farach-Colton et al. (farach-colton:lower-bounds, ). We note a couple things. First, these lower bounds are dramatically different from the lower bound by Willard (willard:loglog, ), where CD is available. Second, Jurdziński and Stachowiak (jurdzinski2002probabilistic, ) demonstrate an algorithm with a matching upper bound of $O(\log n\,\log(1/\epsilon))$ makespan, so long as devices have unique labels, or $n$ is known.
@@
The signal-to-interference-plus-noise ratio (SINR) model has been considered in the context of several fundamental wireless-communication problems (see (moscibroda2006complexity, ; jurdzinski2012distributed, ; jurdzinski2013distributed, ; moses2020deterministic, ; jurdzinski2015cost, ; JurdzinskiR17, )). Informally, under the SINR model, the ratio of (i) the power of an incoming transmission to (ii) the total interference from other devices plus the “background” (or floor) noise, must exceed a threshold value in order for the transmission to succeed. In this setting, Fineman et al. (fineman:contention, ) show how to solve wakeup with $O(\log n+\log R)$ makespan w.h.p., where $R$ is ratio between the longest and shortest link. The authors point out that, in many practical settings, $R$ is at most a polynomial in $n$, which yields an $O(\log n)$ result. Also given is a general lower bound (i.e., not restricted to the class of fair algorithms) of $\Omega(\log n)$.
Finally, a more recent result by Gilbert et al. (gilbert2021contention, ) examines the implications of system-size “predictions”; that is, the algorithm is given access to the probability distribution defined over the possible values of $n$. The authors derive lower and upper bounds for both deterministic and randomized solutions, for the CD and no-CD cases. Interestingly, in addition to new lower bounds, this work recovers some familiar results, such as the lower bound by Willard (willard:loglog, ).
Summary.
For static contention resolution, many randomized algorithms have been proposed, with STB providing one of the strongest results: asymptotically-optimal $O(n)$ makespan with no-CD, no knowledge of $n$, and only $O(\log n)$ channel-access attempts. In contrast, purely deterministic approaches cannot achieve linear makespan, and these results rely on additional assumptions regarding channel feedback, knowledge of $n$, or a unique labeling of participants.
@@
For static wakeup, under each of CD or no-CD, asymptotically matching upper and lower bounds are known. It is interesting to compare these results to those of static contention resolution. In particular, while the makespan for contention resolution is higher, the optimal average per-packet time to success is $\Theta(1)$, which is lower than even the CD case for wakeup.
@@ 3. Dynamic Arrivals
To this end, adversarial queuing theory (AQT) (borodin2001adversarial, ; cruz-one, ; cruz-two, ; andrews2001universal, ) has been used to model packet arrivals times, and many contention-resolution results fit within the AQT framework; that is, they address adversarial models that are parameterized by rates of packet injection and burstiness. Another large body of the contention-resolution literature eschews the parameterization of AQT and instead considers a completely arbitrary schedule for when packets arrive. Here, we treat separately the literature on deterministic and randomized approaches, in Sections 3.2 and 3.3, respectively.
@@ 3.1. AQT and Contention Resolution
Multiple adversarial models appear in the AQT literature; see Borodin et al. (borodin2001adversarial, ) for an in-depth discussion. In the context of contention resolution, there are two prominent models. The leaky-bucket adversary is defined such that, for any $t$ consecutive slots, at most $\rho t+b$ packets may be placed in the queue of any participant (cruz-one, ). For a contiguous segment of rounds, and any positive integer $w$, a window adversary of type $(\rho,w)$ can inject $\rho w$ packets in each contiguous segment of $w$ rounds into any set of stations; $w$ is referred to as a window.
@@
We can now give some intuition for why BEB’ is unstable under an arrival rate of $\Omega(\ln\ln n/\ln n)$; Figure 4 complements our discussion. Consider an interval of $T=\Theta(n\ln n/\ln\ln n)$ slots. In the first slot, a single burst of $n/2$ packets arrives. This drives up the contention to $\Omega(n)$ in this slot. None of these packets is likely to succeed over the next $n/\ln n$ slots; to see this, note that in the $i$-th slot measured from this first slot is sending with probability $1/i$, yielding a contention of $\Omega(\ln n)$ in each of these $n/\ln n$ slots, and so the probability of success is $O(1/{\texttt{poly}({n})})$ by Equation 1.
@@
However, given the burst of packets, the packets in the stream face $n/\ln n$ slots where w.h.p. no successes can occur. Consequently, these packets accumulate in the system and, over these $n/\ln n$ slots, their contribution to contention becomes $\Omega(\ln\ln n/\ln n)\sum_{i=1}^{n/\ln n}1/i=\Omega(\ln\ln n)$ by the sum of a harmonic series. By Equation 1, the probability of success is $O(1/{\texttt{poly}({\ln n})})$, and thus the expected number of slots for a packet to succeed is $\Omega({\texttt{poly}({\ln n})})$.
@@
The remaining approaches in the literature that we survey are deterministic. Chlebus et al. (chlebus2009maximum, ; chlebus2007stability, ) consider adversaries where $\rho=1$; that is, one packet is injected per slot. The main positive result is a deterministic, adaptive protocol that guarantees against leaky-bucket adversaries that $O(n^{2}+b)$ packets are queued in the system at any given time, where $n$ is the number of participants. Here, $n$ is known to all participants and each ID is unique in the range $[1,...,n]$; this allows participants to schedule via passing a token, such that no collisions occur (notably, the positive results in this paper do not require collision detection). This result is shown to be asymptotically optimal in the sense that a leaky-bucket adversary with burstiness $2$ can force $\Omega(n^{2})$ total packets to be queued.
@@
Several impossibility results are also given—holding regardless of whether or not collision detection is available—that illustrate how even minor burstiness can prevent stability. Specifically, no full-sensing protocol can be stable (i) for at least two participants against a leaky-bucket adversary with burstiness $2$, (ii) for at least three participants against a window adversary of burstiness $2$; and (iii) no acknowledgment-based algorithm is stable in a system of at least two participants against the window adversary of burstiness $2$.
@@
Additional impossibility results are made in the context of three properties. First, an algorithm is fair if each packet is transmitted successfully within $O(b/\rho)$ slots from the time it is injected (this is a different notion of fairness than described in Section 2.3). The authors show that no algorithm can be fair and stable for any number of at least two participants against a leaky-bucket adversary of burstiness $2$. Second, an algorithm is withholding the channel if, upon successfully transmitting a packet, the station keeps sending until all of its pending packets are exhausted. No algorithm that withholds the channel can be stable for three participants against a leaky-bucket adversary of burstiness $2$. Third, an algorithm is retaining if after a participant, say $u$, successfully transmits a packet, and if any packets arriving subsequent to this event are injected only at other participants, $u$ eventually empties its queue. No retaining protocol is stable with at least four participants against a window adversary of burstiness $2$. Chlebus et al. show that this implies no fair protocol can also be stable with at least four participants against the window adversary of burstiness $2$.
@@
In contrast to the above, Chlebus et al. (chlebus2006adversarial, ; chlebus:adversarial, ) study deterministic algorithms under the same model as Chlebus et al. (chlebus2009maximum, ; chlebus2007stability, ), but where $\rho<1$. They show that strong stability is impossible for a rate $\omega(1/\log n)$, with or without CD. With CD, the authors give a full-sensing algorithm that is fair and stable for $\rho=O(1/\log n)$, and they show that fairness is not possible for rates that are $\omega(1/\log n)$. When collision detection is not available, the authors give an existence proof for a stable and fair algorithm under a much smaller rate of $O(1/n\log^{2}n)$, and they give an explicit algorithm that achieves these properties for a rate of $O(1/n^{2}\log n)$. Finally, it is shown that no acknowledgement-based algorithm can be stable for a rate exceeding $3/(1+\lg n)$.
@@
Anantharamu et al. (anantharamu2010deterministic, ) and Aldawsari et al. (aldawsari:adversarial, ) design and analyze several deterministic algorithms that employ either a round-robin or token-passing approach for obtaining access to the channel. The authors derive bounds on latency under the leaky-bucket adversary model as a function of $\rho$, for $\rho<1/(2\log n)$, and general $b$. Note that this restriction on $\rho$ should not surprise us, given the negative results on strong stability by Chlebus et al. (chlebus2006adversarial, ; chlebus:adversarial, ) and prior results on the challenges posed by burstiness (chlebus2009maximum, ; chlebus2007stability, ). These algorithms are compared, both analytically and via simulations, against the randomized algorithms, BEB and Quadratic Backoff (QB), where the latter has each successive window size increase as the square of the previous window. The simulation results illustrate the impact of different injection rates, burstiness values, and the number of participants. These findings generally align with the theoretical predictions and, interestingly, the deterministic algorithms compare favorably to BEB and QB, suggesting that randomized approaches may not necessarily be superior. More-recent work by Hradovich et al. (HradovichKK21, ) provides a new average-case analysis of the maximum queue size and latency for several of the same algorithms.
@@
While many deterministic approaches require nodes to have unique IDs, Anantharamu and Chlebus (anantharamu2015broadcasting, ) consider deterministic algorithms where this is not true; furthermore, $n$ is unknown to the algorithm. In trade, they focus on positive results against leaky bucket adversaries that have the additional constraint that they inject at most one packet per round. An algorithm is activation-based if stations that do not have any packets to send ignore any channel feedback. In terms of positive results, the authors provide a non-adaptive, activation-based algorithm that has bounded latency and is stable for an injection rate $\rho<1/3$ with CD, a non-adaptive full-sensing algorithm that has bounded latency and is stable for $\rho\leq 3/8$ with CD, and an adaptive activation-based algorithm that has bounded latency and is stable for injection rates less than $\rho\leq 1/2$ with no-CD.
@@
In terms of negative results, Anantharamu and Chlebus show that no deterministic algorithm can be fair against an adversary that is able to activate at most $2$ stations in any round—called a $2$-activating adversary—with $\lfloor\rho+b\rfloor=1+b\geq 2$. Additionally, the authors show that no deterministic acknowledgment-based algorithm is fair against a $1$-activating adversary for $\rho<1$ and $b\geq 3-2\rho$. This stands in contrast to the positive results in Chlebus et al. (chlebus2006adversarial, ; chlebus:adversarial, ) for $\rho<1$. Finally, Anantharamu and Chlebus prove that no deterministic algorithm with CD can provide bounded latency against a $1$-activating adversary with injection rate at least $3/4$ and $b\geq 1$. It is interesting to compare this result against the algorithm by Chlebus et al. (chlebus2009maximum, ; chlebus2007stability, ) discussed above, which achieves stability, having $O(n^{2}+b)$ total packets queued in the system at any given time (but where there are unique IDs and the number of stations is known).
@@
Anantharamu et al. (anantharamu2009adversarial, ; anantharamu2017adversarial, ) consider deterministic full-sensing and acknowledgment-based algorithms for the case where each participant is subject to an individual injection rate under a window adversary, where the window size $w$ is unknown. There are $n$ participants, each with a unique ID in ${1,...,n}$, and $n$ is known. The authors prove a lower bound on latency of $\Omega(w\max\{1,\log_{w}n\})$ for $w\leq n$, and $\Omega(w+n)$ for $w>n$. Upper bounds are established by a full-sensing algorithm with CD, and by an adaptive algorithm without CD, each achieving a latency of $O(\min(n+w,w\log n))$; Anantharamu et al. also give a full-sensing algorithm without CD that has $O(n+w)$ queue size and $O(nw)$ latency.